# Diffusion MRI with Machine Learning

### v1

@@ II-B Fiber orientation density estimation
Estimation of fiber orientation distribution (FOD), especially in regions with complex fiber configurations, requires high angular resolution diffusion imaging (HARDI) measurements. However, even when such measurements are available, this is a challenging estimation task. Recent studies have highlighted the inherent limitations of FOD estimation based on dMRI data [98, 99]. Nonetheless, the advantages of estimating crossing fiber configurations for applications such as tractography and connectivity analysis has been shown time and time again [100, 101]. As a result, there has been ongoing efforts to improve the accuracy of FOD estimation from dMRI data. The simplest approach to characterizing crossing fibers may be the multi-tensor model [102, 103]. However, these methods require determining the number of tensors in each voxel, which is a difficult model selection problem [104, 105]. At present, spherical deconvolution methods such as constrained spherical deconvolution (CSD) are the most widely used techniques for multi-fiber analysis [106, 107]. These methods consider the dMRI signal in q-space to be the result of the convolution between a spherical point-spread function representing the fiber response function and the FOD. Naturally, they estimate the FOD via deconvolution of the dMRI signal with this response function.
@@
Many machine learning methods have been proposed for FOD estimation in the past decade. Table IV provides a brief listing of some of these methods. We summarize our main observations and conclusions from our study of these methods below.
TABLE IV: A summary of some of the machine learning methods for FOD estimation.
 Reference
Summary
 [108]
@@
This work uses autoencoders to learn FOD priors from high-quality dMRI data. The prior is used to constrain CSD for FOD estimation on test data. Compared with standard estimation techniques, the new method reduces the required number of measurements by a factor of two and can work with a lower diffusion strength (b-value).
[109], [110]
@@
These works are based on rotation- and translation-equivariant convolutional networks. The method is trained in an unsupervised manner by convolving the computed FOD with the tissue response function to predict the dMRI signal and using the error in the predicted signal as optimization loss. The new methods results in lower FOD estimation error and more accurate tractography on in-vivo human data and phantom data.
[111, 112]
@@
These works have developed and validated FOD estimation models using training data with histology-derived FOD ground truth. The models are MLPs with residual connections. Results show that the deep learning methods lead to significantly higher estimation accuracy (quantified in terms of angular correlation coefficient) compared with standard methods such as CSD. On in-vivo human data, the new method shows higher reproducibility.
[113]
@@
A CNN has been used to predict FOD from six diffusion-weighted measurements for neonatal brains. Estimation accuracy is on par with the state of the art on neonatal scans. The method also performs well qualitatively on out-of-distribution clinical datasets of fetuses and newborns.
[114]
@@
The model includes a succession of CNN classifiers. Especially when the number measurements is small, the new method shows superior results compared with the state of the art. It can reconstruct voxels with three crossing fascicles from 10 dMRI measurements, which is a remarkable result.
[115]
@@
This is a lightweight CNN that works on cubic patches of size three voxels. It achieves accurate estimation of FOD with 25 measurements. It estimates crossing fibers better than multi-tissue multi-shell constrained spherical deconvolution (MSMT-CSD [107]).
[116]
@@
This method, named Fiber Orientation Reconstruction guided by a Deep Network (FORDN), is based on overcomplete dictionaries and MLP networks. A coarse dictionary is first used to represent the signal. In the second stage, a larger dictionary is used to compute a finer FOD that is close (in an $\ell_{1}$ sense) to the coarse FOD computed in the first step. Compared with methods that are based on sparse reconstruction or deep learning methods, this method is more accurate especially in voxels with two and three crossing fibers.
[117]
@@
This work uses a voxel-wise MLP that is trained on simulated or real data. Extensive experiment show that this method is superior to a range of conventional FOD estimation methods when the number of measurements is small.
[118, 119]
@@
These works apply 2D and 3D CNNs on data from individual voxels or small patches to compute the number of fascicles and the complete FOD. Experimental results on HCP data are reported that show that the deep learning methods can estimate the number of major fascicles and the FOD in voxels with complex fiber configurations. The deep learning methods can perform these computations with as few as 15 measurements and they are more accurate than CSD.
[120]
@@
To enable the method to work with different q-space shells, the model has three input heads for three common b-values (1000, 2000, and 3000). Either one of the shells or any combination thereof can be supplied at test time. The architecture itself is a CNN.
[121]
@@
A novel feature vector is proposed based on the decay of the diffusion signal as a function of orientation. Using this hand-crafted feature vector as input, an MLP is trained to estimate the “angle to the closest fascicle” for a large set of directions on the sphere. This information is used to infer the number and orientation of the major fascicles or to approximate the complete FOD.
[122]
The main novelty of this work consists in a loss term to encourage scan-rescan reproducibility of the FOD estimates. The method achieves better scan-rescan consistency on external data in terms of FOD estimation as well as structural connectivity metrics.
[123]
@@
The network is inspired by a reformulation of CSD, hence it is presented as a “model-driven deep learning method”. The reformulation is turned into an iterative optimization method that is unfolded and implemented as a deep neural network. The loss function has a $\ell_{2}$ term for the predicted FOD and a cross entropy term for the predicted number of fixels (i.e., major peaks).
@@
  • Prior to the surge of deep learning, several studies used classical machine learning methods for FOD estimation [124, 125]. One study computed priors from a population of training subjects and employed the learned prior in a maximum a posteriori framework to improve the estimation accuracy on test data [125].
@@
  • In more recent years, a number of studies have trained neural networks to estimate the FOD from dMRI data [126, 119]. These studies are different in terms of the neural network architecture, training strategy, and evaluation metrics. However, most share a similar claim that a deep neural network can estimate FODs more accurately than spherical deconvolution methods.
@@
  • Methodologically, most deep learning studies have followed a straight estimation approach, where the input dMRI measurements are mapped to the target FOD. The input signal as well as the target FOD are usually expressed as either functions on a discrete spherical grid or using spherical harmonics. However, there have been many exceptions, three of which we briefly describe here. One study has proposed to use an auto-encoder to learn a model of plausible FOD shapes from a high-quality training dataset [108]. This prior is then used to regularize the CSD method to obtain more accurate predictions than the standard CSD. Results show that this technique outperforms conventional methods when measurements are few or the diffusion strength (i.e., the b-value) is low. Koppers and Merhof [114] use a succession of two CNN classifiers, where the first CNN estimates the number (either 1, 2, or 3) of fibers in the voxels and the second CNN estimates the fiber orientations. Ye and Prince [116] used deep neural networks to estimate the FOD by solving sparse estimation problems in dictionaries. A coarse dictionary was used in the first stage to compute an initial estimate, which was then refined using a finer dictionary. Bartlett et al. start by reformulating the CSD equation to derive an iterative optimization algorithm for FOD estimation [123]. They propose to solve this problem using a deep neural network. In addition to the $\ell_{2}$ term of the FOD prediction error, they introduce a cross entropy loss term to encourage correct prediction of the number of major fascicles. Their experiments show that their method performs better than MSMT-CSD in estimating FOD and its peaks.
  • The majority of studies have used standard estimation methods such as CSD to generate target FODs for training and validation data [119, 120]. Typically, these works have aimed at matching the standard technique while using fewer measurements as input. It has been claimed that neural networks can use 6-20 measurements from a single shell to estimate the FOD estimated from multi-shell HARDI data computed with standard methods. Therefore, the validation has been based on comparison with standard methods. This is unlike the common validation approach for FOD estimation techniques, which has been primarily based on simulation [127]. However, with this approach, the trained model inevitably inherits some of the shortcomings of the standard method that is used to compute the target FOD [109, 98, 99]. Moreover, the assessments can only tell us how close the new machine learning method is to the standard technique.
  • A few important studies have used histology data for training and/or validation. Nath et al. used ground truth histology FOD data from squirrel monkey brains, registered to dMRI data, to assess and compare a deep learning technique and CSD [111, 112]. Their experimental results showed that there was additional information in the dMRI measurements that CSD failed to utilize, and that a deep learning method was capable of using this extra information for FOD estimation. The deep learning model, which was a voxel-wise MLP, outperformed CSD in terms of estimation accuracy using histology as ground truth. Further experiments with in-vivo human scans from the HCP dataset showed better scan-rescan reproducibility of the deep learning method, which the authors interpret as evidence that the method can be used in clinical applications. Other studies [109, 110] have proposed novel methods for sidestepping the need for an FOD ground truth. They compute the convolution of their estimated FOD with the fiber response function to compute the corresponding dMRI signal and compute the training loss as the difference between this predicted signal and actual measurements.
@@
  • Unlike the scalar micro-structural indices, FOD is a function of angle and, hence, it is rotation-sensitive. Elaldi et al. argue that the standard convolutional layers (which provide equivariance to planar translation) are insufficient for spherical signals in dMRI, where rotation equivariance is additionally needed [109]. Instead, they employ rotation-equivariant graph convolutions proposed in [128]. Furthermore, they opt for spherical harmonics of degree 20 (as opposed to degree 6 or 8 in standard methods [106]) to enable reconstruction of nearby FOD peaks. They show that, compared with the CSD baseline, their neural network achieves superior FOD estimation and better performance on downstream tasks such as tractography on phantom data and human brain scans. They have further extended this work by developing neural network layers that are equivariant to translations and grid reflections (as are standard FCNs) as well as to voxel and grid rotations [110]. They argue that these extra equivariance properties are essential to give the network the necessary spatio-spherical inductive biases to learn to compute the FOD from data. Their results show more accurate FOD estimation, tractography, and brain tissue segmentation in dMRI compared with a range of state of the art methods. We review related neural network architectures in Section III-C.
@@
  • Instead of estimating the complete FOD, several studies have addressed the less ambitious but still challenging problem of determining the number of major fascicles in each voxel. Schultz used support vector regression to estimate the number of major fascicles and trained the model with simulated data [124]. Evaluations on simulated and real brain data showed that this learning method was able to determine the number of major fascicles more accurately than CSD. Another study [121] developed novel feature vectors to characterize the decay of the diffusion signal as a function of orientation. The feature vectors were then used by an MLP to compute the angle to the closest fascicle. This information was further processed via smoothing and local minimum extraction to determine the number and orientation of major fascicles in the voxel.
II-C Tractography and structural connectivity
@@
Tractography is one of the most common and most controversial computations enabled by dMRI [129, 130, 131]. Early tractography methods relied on the diffusion tensor model for computing the local fiber orientations [132]. Over the past two decades, more advanced streamline tracing methods have been developed including anatomically constrained tractography [133], global tractography algorithms [134], ensemble tractography [135], augmented tractography [136], and micro-structure-informed tractography filtering [137]. These methods aim to utilize anatomical context information or the correspondence between tractography streamlines and the local dMRI signal to overcome the limitations of conventional tractography techniques. However, tractography is intrinsically ill-posed and suffers from high false positive and false negative rates [19, 20]. Some of the main sources of error, unreliability, and ambiguity in tractography include [130, 138, 139, 140]: (1) Difficulty of modeling crossing fibers, (2) Ambiguity of streamline tracing in voxels with crossing or kissing fibers, (3) Presence of bottleneck regions, which refer to regions of the white-matter where several bundles pass in the same direction before spreading, and (4) Anatomical biases such as gyral bias and termination bias. Standard tractography methods often produce inaccurate results that can influence tract-specific analysis and quantitative structural connectivity assessment [141, 130, 138].
@@
Machine learning methods may offer a framework for developing better tractography algorithms. They are model-free and allow for a seamless integration of anatomical priors and other sources of information that may be useful for tractography. As a result, they have the potential to have lower false positive rates, to be more robust to measurement noise, and to overcome the position and shape biases of classical tractography algorithms [22, 142, 143]. Although various methods such as self-organizing maps [144, 145] and random forests [142, 39] have been successfully used for tractography, machine learning-based tractography techniques have increasingly relied on neural networks [22, 142, 146, 147]. Table V lists some of the works that have developed machine learning methods for streamline tractography. We summarize the main technical aspects of these methods and their reported results below.
@@
TABLE V: A summary of selected machine learning methods for streamline tractography.
 Reference
Model (D: deterministic, P: probabilistic)
Model input
Main findings
[39, 142]
random forests (D & P)
@@
Raw diffusion signal in a small neighborhood of the current voxel & direction of the three preceding streamline steps.
Higher sensitivity and specificity and lower angular error than 12 state of the art methods.
[150]
GRU (D)
Raw diffusion signal in the current voxel.
@@
Better tract coverage and lower false positive rate than classical methods.
[148]
RNN (D & P)
Raw diffusion signal in the current voxel.
@@
Lower percentage of invalid connections and non-connections, higher rate of valid bundles, and higher bundle coverage than a range of learning and classical methods.
[152]
MLP (D)
Raw diffusion signal & direction of the two preceding streamline steps.
@@
This work reports extensive experiments to determine the optimal input and network architectures.
[42]
MLP (D)
Raw diffusion signal in a $3^{3}$-voxel neighborhood & direction of the two preceding streamline steps.
@@
Better results than standard tractography methods when trained on small datasets. The method is robust to noisy training data.
[44]
convolutional-recurrent neural network
Anatomical MRI and streamline memory.
@@
The variability in tractography and structural connectivity measures computed based on T1w MRI is similar to that for standard dMRI-based results.
[151]
deep reinforcement learning (D & P)
FOD at the current voxel and its six neighbors & direction of the past four streamline steps.
@@
A reinforcement learning approach can be competitive with supervise learning methods while also being more generalizable to unseen or out-of-distribution data.
[147]
GRU (D)
dMRI signal.
@@
Superior to classical methods in terms of true positive rate and better bundle volume coverage than existing probabilistic techniques.
[149]
MLP (P)
Direction of the three last streamline steps & FOD, represented as SH of order 4, in a voxel neighborhood of size 3 voxels.
@@
The method achieves competitive results compared with a range of conventional and machine learning methods on independent synthetic test data.
  • Machine learning-based methods have been reported to be effective in reconstructing complex streamlines with low false positive rates. They have the potential to reduce the systematic errors and biases that are prevalent in classical tractography methods and to improve generalizability to unseen data [39, 22]. Several reasons have been put forward to explain the success of machine learning methods. First, they avoid making rigid, ad-hoc, and possibly sub-optimal modeling assumptions or streamline propagation and stopping rules that are common in standard methods [39, 22]. Some of the proposed machine learning methods use the dMRI data (rather than diffusion tensor or FOD estimates) as input, thereby sidestepping the unavoidable errors in fiber orientation estimation [39, 148, 42]. Moreover, they have the potential to learn tissue probabilities from dMRI data, rather than relying on the information provided by a registered anatomical MRI or unreliable parameters such as FA. Machine learning methods can also learn the noise and artifacts from data, instead of imposing simplified noise models and assuming artifact-free scans [39]. Furthermore, these methods can incorporate estimation uncertainty into streamline tractography in a systematic way [149]. Machine learning techniques are also by far more flexible in incorporating non-local information, such as anatomical context, which can greatly improve tractography results [22]. An important piece of information that can be easily incorporated into most machine learning models is the direction of preceding streamline tracing steps. The direction of the last step indicates the streamline slope and the direction of the last two steps indicate its curvature. This information has been useful in improving tractography accuracy [150]. Incorporation of the direction of preceding steps can also improve the tractography in voxels with crossing fibers and bottleneck regions [151, 150].

### v2

@@ IV-E Fiber orientation density estimation
Estimation of fiber orientation distribution (FOD), especially in regions with complex fiber configurations, requires high angular resolution diffusion imaging (HARDI) measurements. However, even when such measurements are available, this is a challenging estimation task. Recent studies have highlighted the inherent limitations of FOD estimation based on dMRI data [183, 184]. Nonetheless, the advantages of estimating crossing fiber configurations for applications such as tractography and connectivity analysis has been shown time and time again [185, 186]. As a result, there have been ongoing efforts to improve the accuracy of FOD estimation. The simplest approach to characterizing crossing fibers may be the multi-tensor model [187, 188]. However, these methods require determination of the number of tensors in each voxel, which is a difficult model selection problem [189, 190]. At present, spherical deconvolution methods such as constrained spherical deconvolution (CSD) are the most widely used techniques for assessing crossing tracts [191, 192]. These methods consider the dMRI signal in q-space to be the result of the convolution between a spherical point-spread function representing the fiber response function and the FOD. Naturally, they estimate the FOD via deconvolution of the dMRI signal with this response function.
@@
Many machine learning methods have been proposed for FOD estimation. Table V provides a brief listing of some of these methods. We summarize our main observations and conclusions from our study of these works below.
TABLE V: A summary of some of the machine learning methods for FOD estimation.
  Method Summary of methodology and results
@@
  Autoencoder-based FOD regularization [193] This work uses autoencoders to learn FOD priors from high-quality dMRI data. The prior is used to constrain CSD for FOD estimation on test data. Compared with standard estimation techniques, the new method reduces the required number of measurements by a factor of two and can work with a lower diffusion strength (b-value). The method is tested on only one subject.
@@
Equivariant networks [194], [195] These works are based on rotation- and translation-equivariant convolutional networks. Training is performed in an unsupervised manner by convolving the computed FOD with the tissue response function to predict the dMRI signal and using the error in the predicted signal as optimization loss. The new methods results in lower FOD estimation error and more accurate tractography on in-vivo human data and phantom data.
@@
MLP trained with histology ground truth [196] These works have developed and validated FOD estimation models using training data with histology-derived FOD ground truth. The models are MLPs with residual connections. Results show that the deep learning methods lead to higher estimation accuracy (quantified in terms of angular correlation coefficient) compared with standard methods such as CSD. On in-vivo human data, the new method shows higher reproducibility.
@@
CNN method for fetal and neonatal brains [197] A CNN has been used to predict FOD from six diffusion-weighted measurements for neonatal brains. Estimation accuracy is on par with the state of the art on neonatal scans. The method also performs well qualitatively on out-of-distribution clinical datasets of fetuses and newborns.
@@
CNN classifiers [198] The method, which is based on CNN classifiers, shows better results than the state of the art, especially when the number of measurements is small. It can reconstruct voxels with three crossing fascicles from 10 dMRI measurements, but the method is only tested on synthetic data.
@@
Lightweight CNN [199] The model is a CNN that works on cubic patches of size three voxels. It achieves accurate estimation of FOD with 25 measurements. It estimates crossing fibers better than multi-tissue multi-shell constrained spherical deconvolution (MSMT-CSD [192]).
@@
FORDN [200] This method, named Fiber Orientation Reconstruction guided by a Deep Network (FORDN), is based on overcomplete dictionaries and MLP networks. A coarse dictionary is first used to represent the signal. In the second stage, a larger dictionary is used to compute a finer FOD that is close (in an $\ell_{1}$ sense) to the coarse FOD computed in the first step. Compared with methods that are based on sparse reconstruction or deep learning methods, this method is more accurate especially in voxels with two and three crossing fibers. However, the method is tested on phantom data and in-vivo scan of one human subject.
@@
Voxel-wise MLP [201] This work uses a voxel-wise MLP that is trained on simulated or real data. Extensive experiment show that this method is superior to a range of conventional FOD estimation methods when the number of measurements is small.
@@
Voxel-wise or small-patch CNNs [202, 203] These works apply 2D and 3D CNNs on data from individual voxels or small patches to compute the number of fascicles and the complete FOD. Experiments on HCP data show the deep learning methods can estimate the number of major fascicles and the FOD in voxels with complex fiber configurations with as few as 15 measurements and they are more accurate than CSD.
@@
Method for heterogeneous multi-shell data [204] To enable the method to work with different q-space shells, the model has three input heads for three common b-values (1000, 2000, and 3000). Either one of the shells or any combination thereof can be supplied at test time. The architecture itself is a CNN. Only HCP data is used to validate the method.
@@
MLP applied on dMRI signal decay features [205] A novel feature vector is proposed based on the decay of the diffusion signal as a function of orientation. Using this hand-crafted feature vector as input, an MLP is trained to estimate the “angle to the closest fascicle” for a large set of directions on the sphere. This information is used to infer the number and orientation of the major fascicles or to approximate the complete FOD.
@@
Spherical deconvolution network [206] The network is inspired by a reformulation of CSD, hence it is presented as a “model-driven deep learning method”. The reformulation is turned into an iterative optimization method that is unfolded and implemented as a deep neural network. The loss function has an $\ell_{2}$ term for the predicted FOD and a cross entropy term for the predicted number of fixels (i.e., major peaks). The study has used HCP data only.
IV-E1 Method design
@@
Prior to the surge of deep learning, a few studies used classical machine learning methods for FOD estimation [207, 208]. One study computed priors from a population of training subjects and employed the learned prior in a maximum a posteriori framework to improve the estimation accuracy on test data [208].
@@
In more recent years, a number of studies have trained neural networks to estimate the FOD [209, 203]. These studies are different in terms of the neural network architecture, training strategy, and evaluation metrics. However, most share a similar claim that a deep neural network can estimate FODs more accurately than spherical deconvolution methods. Methodologically, most deep learning studies have followed a straight estimation approach, where the input dMRI measurements are mapped to the target FOD. The input signal as well as the target FOD are usually expressed either as functions on a discrete spherical grid or using spherical harmonics. However, there have been many exceptions, three of which we briefly describe below.
@@
One study has proposed to use an auto-encoder to learn a model of plausible FOD shapes from a high-quality training dataset [193]. This prior is then used to regularize the CSD method to obtain more accurate predictions than the standard CSD. Results show that this technique outperforms conventional methods when measurements are few or the diffusion strength (i.e., the b-value) is low. Koppers and Merhof [198] use a succession of two CNN classifiers, where the first CNN estimates the number (either 1, 2, or 3) of fibers in the voxel and the second CNN estimates the fiber orientations. Ye and Prince [200] used deep neural networks to estimate the FOD by solving sparse estimation problems in dictionaries. A coarse dictionary was used in the first stage to compute an initial estimate, which was then refined using a finer dictionary. Bartlett et al. start by reformulating the CSD equation to derive an iterative optimization algorithm for FOD estimation [206]. They propose to solve this problem using a deep neural network. In addition to an $\ell_{2}$ loss for FOD prediction error, they introduce a cross entropy loss to encourage correct prediction of the number of major fascicles. Their experiments show that their method performs better than MSMT-CSD in estimating FOD and its peaks.
@@
Unlike the scalar microstructural indices, FOD is a function of angle and, hence, it is rotation-sensitive. Elaldi et al. argue that the standard convolutional layers (which provide equivariance to planar translation) are insufficient for spherical signals in dMRI, where rotation equivariance is additionally needed [194]. Instead, they employ rotation-equivariant graph convolutions proposed in [210]. Furthermore, they opt for spherical harmonics of degree 20 (as opposed to degree 6 or 8 in standard methods [191]) to enable reconstruction of nearby FOD peaks. They show that, compared with the CSD baseline, their neural network achieves superior FOD estimation and better performance on downstream tasks such as tractography on phantom data and human brain scans. They have further extended this work by developing neural network layers that are equivariant to translations and grid reflections (as are standard FCNs) as well as to voxel and grid rotations [195]. They argue that these extra equivariance properties are needed to give the network the necessary spatio-spherical inductive biases to effectively learn to compute the FOD from data. Their results show more accurate FOD estimation, tractography, and brain tissue segmentation in dMRI compared with a range of state of the art methods. Related neural network architectures are reviewed in more detail in Section V-C.
IV-E2 Training data
The majority of studies have used standard estimation methods such as CSD to generate target FODs for training and validation data [203, 204]. Typically, these works have aimed at matching the standard technique while using fewer measurements as input. It has been claimed that neural networks can use 6-20 measurements from a single shell to estimate the FOD estimated from multi-shell HARDI data computed with standard methods. Therefore, the validation has been based on comparison with standard methods. This is unlike the common validation approach for FOD estimation techniques, which has been primarily based on simulation [211]. However, with this approach, the trained model will inevitably inherit some of the shortcomings of the standard method that is used to generate the target FOD [194, 183, 184]. Moreover, the assessments can only tell us how close the new machine learning method is to the standard technique.
A few important studies have used histology data for training and/or validation. Nath et al. used ground truth histology FOD data from squirrel monkey brains, registered to dMRI data, to assess and compare a deep learning technique and CSD [196]. Their experimental results showed that there was additional information in the dMRI measurements that CSD failed to utilize, and that a deep learning method was capable of using that extra information for FOD estimation. The deep learning model, which was a voxel-wise MLP, outperformed CSD in terms of estimation accuracy using histology as ground truth. Further experiments with in-vivo human scans from the HCP dataset showed better scan-rescan reproducibility of the deep learning method, which the authors interpreted as evidence that the method could be used in clinical applications. Other studies [194, 195] have proposed novel methods for sidestepping the need for an FOD ground truth. They compute the convolution of the estimated FOD with the fiber response function to obtain the corresponding dMRI signal and use the difference between this predicted signal and actual measurements as the optimization loss for model training.
IV-E3 Prediction of the number of main fascicles
@@
Instead of estimating the complete FOD, several studies have addressed the less ambitious but still challenging problem of determining the number of major fascicles in each voxel. Schultz trained a support vector regression model to estimate the number of major fascicles using simulated training data [207]. Evaluations on simulated and real brain data showed that this method determined the number of major fascicles more accurately than CSD. Another study [205] devised novel feature vectors to characterize the decay of the diffusion signal as a function of orientation. The feature vectors were then used by an MLP to compute the angle to the closest fascicle. This information was further processed via smoothing and local minimum extraction to determine the number and orientation of major fascicles in the voxel. Comparisons with several classical methods showed that this machine learning technique was more accurate.
IV-F Tractography and structural connectivity
@@
Tractography algorithms build on local fiber orientations to compute virtual streamlines that connect different brain regions [212]. Tractography has important applications, most prominently delineation of specific white matter tracts and quantitative structural connectivity [213, 214]. It is one of the most common, challenging, and controversial computations enabled by dMRI [215, 216, 217]. Early tractography methods relied on the diffusion tensor model for computing the local fiber orientations [218]. Over the past two decades, more advanced streamline tracing methods have been developed including anatomically constrained tractography [219], global tractography algorithms [220], ensemble tractography [221], augmented tractography [222], and microstructure-informed tractography filtering [223]. These methods aim to utilize anatomical context information or the correspondence between tractography streamlines and the local dMRI signal to overcome the limitations of conventional tractography techniques.
However, tractography is intrinsically ill-posed and suffers from high false positive and false negative rates [37, 38]. Some of the main sources of error, unreliability, and ambiguity in tractography include [216, 213, 224, 225]: (1) Difficulty of modeling crossing fibers, (2) Ambiguity of streamline tracing in voxels with crossing or kissing fibers, (3) Presence of bottleneck regions, where several tract bundles trace the same voxels in the same direction, and (4) Anatomical biases such as the gyral bias and termination bias. Standard tractography methods often produce inaccurate results that can influence tract-specific analysis and structural connectivity assessment [226, 216, 213].
@@
Machine learning may offer a framework for developing better tractography algorithms. They are model-free and allow for a seamless integration of anatomical priors and other sources of information that may be useful for tractography. Some of these information, such as the direction of previous tractography steps, have been used in conventional tractography methods in the past [227, 228]. Nonetheless, machine learning methods offer higher flexibility in integrating these various inputs into a unified model and to optimize the model to reduce the tractography error with respect to all inputs jointly. As a result, they have the potential to have lower false positive rates, to be more robust to measurement noise, and to overcome the inherent biases of classical tractography algorithms [18, 229, 230]. Although various methods such as self-organizing maps [231] and random forests [229] have been successfully used for tractography, machine learning-based tractography techniques have increasingly relied on neural networks [18, 229, 232, 233]. Table VI lists some of the works that have developed machine learning methods for streamline tractography. We summarize the main technical aspects of these methods and their reported results below.
@@
TABLE VI: A summary of selected machine learning methods for streamline tractography.
  Reference Model (D: deterministic, P: probabilistic) Model input Main findings
@@
  Fiber tracking with a random forest classifier [229] Random forests (D & P) Raw diffusion signal in a small neighborhood of the current voxel & direction of the three preceding streamline steps. Higher sensitivity and specificity and lower angular error than 12 state of the art methods.
@@
Learn to Track [234] GRU (D) Raw diffusion signal in the current voxel. Better tract coverage and lower false positive rate than classical methods.
@@
DeepTract [235] RNN (D & P) Raw diffusion signal in the current voxel. Lower percentage of invalid connections and non-connections, higher rate of valid bundles, and higher bundle coverage than a range of learning and classical methods.
@@
Jörgens et al. [236] MLP (D) Raw diffusion signal & direction of the two preceding streamline steps. This work reports extensive experiments to determine the optimal input and network architectures.
@@
Wegmayr et al. [59] MLP (D) Raw diffusion signal in a $3^{3}$-voxel neighborhood & direction of the two preceding streamline steps. Better results than standard tractography methods when trained on small datasets. The method is robust to noisy training data.
@@
Cai et al. [61] Convolutional-recurrent neural network Anatomical MRI and streamline memory. The variability in tractography and structural connectivity measures computed based on T1-weighted MRI is similar to that for standard dMRI-based results.
@@
Track-to-Learn [237] Deep reinforcement learning (D & P) FOD at the current voxel and its six neighbors & direction of the past four streamline steps. A reinforcement learning approach can be competitive with supervise learning methods while also being more generalizable to unseen or out-of-distribution data.
@@
Bundle-Wise Deep Tracker [233] GRU (D) dMRI signal. Superior to classical methods in terms of true positive rate and better bundle volume coverage than existing probabilistic techniques.
@@
Entrack [238] MLP (P) Direction of the three last streamline steps & FOD, represented as SH of order 4, in a voxel neighborhood of size 3 voxels. The method achieves competitive results compared with a range of conventional and machine learning methods on independent synthetic test data.
IV-F1 Method design